OpenStudy (anonymous):

Please help! x^2+12x+20=0

OpenStudy (anonymous):

Solve for x by completing the square and factoring, I am not completely sure how to do this, can someone help me?

OpenStudy (anonymous):

6

OpenStudy (idealist10):

Factor, x^2+12x+20=0 (x+10)(x+2)=0 x+10=0, x+2=0 x=-10, -2. Answer: -10, -2. The key here is to find 2 numbers that when they multiply, you get the third term and when you add them, you get the second term. So for this problem, these 2 numbers are 10 and 2. Because 10 times 2 gives you 20 and 10+2=12.

OpenStudy (anonymous):

Ok! Now I understand, thank you so much Idealist10!!!!

OpenStudy (idealist10):

You're welcome!